A body of mass 60 kg is dragged with just enough force to start moving on a rough surface with coefficients of static and kinetic friction 0.5 and 0.4 respectively.On applying the same force what is the acceleration?
a..98 m/s2 b.9.8 m/s2 c. .54 m/s2 d 5.292m/s2

Grade:12

1 Answers

Aakash Anuj
13 years ago

See the maximum static friction that can act on the body is Umg = 0.5*60*9.8 = 294 N

now since the force is just above 300 N, the body will translate and the type of friction now will be kinetic instead of static. Hence the net acceleration is 294-0.4*60*9.8/60 = 0.98 m/s^2 Ans.